OpenStudy (anonymous):

Help me please. best answer gets 12 rewards

OpenStudy (tester97):

what do you need help with?

OpenStudy (anonymous):

Hold on

OpenStudy (anonymous):

What is the 25th term of the geometric sequence that starts with .01 and has a ratio of 2?

OpenStudy (anonymous):

For a geometric sequence, the nth term can be given as an = a1 r^(n-1) where an = nth term a1 = first term r = common ratio so just substitute in the values a25 = 0.01 * 2^(25-1) = 167,772

OpenStudy (anonymous):

Okay, Thanks!!!1

OpenStudy (anonymous):

My teacher said something about nth but I was still confused.

OpenStudy (tester97):

yea i was about to say that

OpenStudy (tester97):

good job non the less ;)

OpenStudy (anonymous):

"nth term" is just like "xth term". x (or n) is the unknown number